# EBC-ZIP: Improving Blockwise Crowd Counting with Zero-Inflated Poisson Regression

### v1

@@ III-C Extending EBC to EBC-ZIP
Suffix Backbone Size (M) FLOPS (G)
-P (Pico) MobileNetV4-Small-0.5×0.5$\times$0.5 × 0.81 6.46
-N (Nano) MobileNetV4-Small 3.36 24.73
-T (Tiny) MobileNetV4-Medium 10.64 61.62
-S (Small) MobileCLIP-S1 33.72 242.68
-B (Base) CLIP-ConvNeXt-Base 105.60 800.99
TABLE I: EBC-ZIP model variants with corresponding backbones, parameter sizes, and computational complexity measured on $1920\times 1080$ resolution.
Under the assumption that blockwise counts are conditionally independent, the negative log-likelihood (NLL) loss of the zero-inflated Poisson distribution is given by:
$\mathcal{L}_{\mathrm{NLL}}=$ $-\log\prod_{i=1}^{h}\prod_{j=1}^{w}P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$
$=$ $-\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (10)
where $P(\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ is defined by the ZIP distribution in Eq. (4), and $\boldsymbol{\lambda}$ and $\boldsymbol{\pi}$ are given by (7) and (9), respectively. It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss Eq. (10). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
@@
$\boldsymbol{Y}^{*}_{i,j}=\mathbb{E}[\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j}].$ (11)
Using Eq. (5), $\boldsymbol{Y}^{*}$ can be expressed as
$\boldsymbol{Y}^{*}=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda},$ (12)
@@
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of EBC-ZIP with different backbones, including MobileNetV4 [21], MobileCLIP [22] and OpenCLIP’s ConvNeXt-Base [24, 23] These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table I. All variants share the same head design and differ only in backbone complexity. This modularity allows EBC-ZIP to scale across computational budgets.
Regression Component MAE RMSE NAE
MAE 6.69 10.45 5.52%
DMCount [11] 6.62 9.83 5.30%
Poisson NLL 6.57 9.86 5.47%
ZIP NLL (ours) 6.36 10.32 5.00%
TABLE II: Comparison of ZIP NLL with other choices in the total loss function Eq. (15) on ShanghaiTech B. Other loss terms (i.e., $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$) remain unchanged. Results show that our ZIP NLL achieves the lowest MAE and NAE.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU-Crowd (Val)
MAE RMSE MAE RMSE MAE RMSE MAE RMSE
BL [10] 62.8 101.8 7.7 12.7 88.7 154.8 93.6 470.3
DMCount [11] 59.7 95.7 7.4 11.8 85.6 148.3 70.5 357.6
EBC [15] 62.3 98.9 7.0 10.9 77.2 130.4 39.6 95.8
EBC-ZIP (ours) 58.6 96.0 6.3 10.3 73.0 126.1 38.3 90.3
Improvement over EBC 5.9% 2.9% 10.0% 5.5% 5.4% 3.2% 3.2% 5.7%
TABLE III: Comparison of EBC-ZIP with other frameworks on ShanghaiTech A & B, UCF-QNRF and NWPU-Crowd. All frameworks use the VGG19-based backbone. Results demonstrate that our framework EBC-ZIP significantly outperforms other methods, achieving the lowest MAE and RMSE on all four benchmarks. Our EBC-ZIP also consistently improves the EBC framework, at most by 10.0%. The best results are highlighted in bold font, and the second best results are underscored.
IV Experiments
Datasets. Following [15], we evaluate all EBC-ZIP variants listed in Table I on four widely used crowd counting benchmarks: ShanghaiTech A & B [5], UCF-QNRF [6], and NWPU-Crowd [7]. ShanghaiTech A contains 300 training images and 182 test images, with highly congested scenes and an average of 501 people per image. ShanghaiTech B, the only dataset collected from surveillance viewpoints, includes 400 training and 316 test images, featuring relatively sparse crowds. UCF-QNRF comprises 1,201 training images and 334 test images with extremely dense crowds, averaging 815 people per image. NWPU-Crowd is currently the largest high-resolution crowd counting dataset, containing 3,109 training images, 500 validation images, and 1,500 test images, with an average count of 418. The test set annotations of NWPU-Crowd are not publicly released. To obtain evaluation results on this split, predicted counts must be submitted to the official evaluation server. This protocol helps maintain a fair comparison between methods by preventing result tweaking using test labels.
@@ IV Experiments
Setup. Our training configuration largely follows [15]. We first preprocess all datasets so that the minimum edge length is no less than 448. We limit the maximum edge length of UCF-QNRF to be 2048 and that of NWPU-Crowd to be 3072. On ShanghaiTech A & B and NWPU-Crowd, we use a block size of 16 while on UCF-QNRF we use a block size of 32. We apply a combination of data augmentation techniques, including random resized cropping, horizontal flipping, color jittering, and salt-and-pepper noise injection. For ShanghaiTech A and B, which consist of lower-resolution images, the crop size is set to $448\times 448$. For the higher-resolution datasets UCF-QNRF and NWPU-Crowd, we use a larger crop size of $672\times 672$. All models are optimized using the Adam optimizer with a weight decay of $41e-$ 4. The learning rate is linearly warmed up from $51e-$ 5 to $41e-$ 4 over the first 25 epochs, followed by cosine annealing with parameters $T_{0}=5$ and $T_{\mathrm{mult}}=2$. Each model variant is trained on each dataset for a total of 1,300 epochs with a batch size of 8. All experiments are conducted on a single NVIDIA RTX 5000 GPU.
@@
where $M$ denotes the total number of test images, $c_{m}$ is the ground-truth count for image $m$, and $c_{m}^{*}$ is the predicted count, computed via Eq. (13). The set $\mathcal{M}+\subseteq{1,2,\dots,M}$ includes only the indices for which $c_{m}>0$, and $|\mathcal{M}+|$ denotes its cardinality. The NAE metric excludes empty scenes to avoid division by zero and to better reflect relative error in populated scenes.
IV-A Effectiveness of EBC-ZIP
In this subsection, we empirically validate the effectiveness of our proposed EBC-ZIP framework by comparing it against several strong baselines, including BL [10], DMCount [11], and EBC [15]. For a fair comparison, we adopt a consistent encoder-decoder architecture based on VGG19 [32]. Before proceeding to the comparison, we first determine the optimal value of the weighting parameter $\omega$ in Eq.(15), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig.5. As illustrated, the model achieves the best MAE and RMSE when $\omega=0.75$, and the lowest NAE when $\omega=1.00$. In our subsequent comparisons, we fix $\omega=1.00$ to prioritize NAE, due to its robustness to varying crowd densities.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function (15) with several commonly used alternatives, including MAE, DMCount [11], and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table II. As shown, our ZIP NLL achieves the lowest MAE (6.36) and NAE (5.00%), and performs competitively in RMSE (10.32), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
Finally, we compare EBC-ZIP framework with several existing methods on the four benchmarks, shown in Table III. Our approach sets a new state-of-the-art on ShanghaiTech B, UCF-QNRF and NWPU-Crowd. It also achieves the lowest MAE and the second lowest RMSE on ShanghaiTech A. Notably, when compared directly with its EBC baseline, our EBC-ZIP model demonstrates significant performance gains across all datasets, reducing the MAE by a margin of 3.2% to 10.0%. hese improvements validate the efficacy of incorporating zero-inflated Poisson regression for crowd counting.
Method Venue ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
@@ IV-A Effectiveness of EBC-ZIP
ChfL [33] CVPR’22 57.5 94.3 6.9 11.0 80.3 137.6 76.8 343.0
MAN [34] CVPR’22 56.8 90.3 - - 77.3 131.5 76.5 323.0
CLTR [35] ECCV’22 56.9 95.2 6.5 10.6 85.8 141.3 61.9 246.3
CrowdHat [36] CVPR’23 51.2 81.9 5.7 9.4 75.1 126.7 - -
STEERER [13] ICCV’23 54.5 86.9 5.8 8.5 74.3 128.3 54.3 238.3
PET [37] ICCV’23 49.3 78.8 6.2 9.7 79.5 144.3 58.5 238.0
APGCC [38] ECCV’24 48.8 76.7 5.6 8.7 80.1 133.6 - -
Gramformer [39] AAAI’24 54.7 87.1 - - 76.7 129.5 - -
mPrompt [18] CVPR’24 52.5 88.9 5.8 9.6 72.2 133.1 50.2 219.0
CLIP-EBC [15] ICME’25 52.5 83.2 6.0 10.1 80.3 136.6 36.6 81.7
EBC-ZIP-B (ours) - 47.81 75.04 5.51 8.63 69.46 121.88 28.26 64.84
@@
TABLE IV: Comparison of our model EBC-ZIP-B with state-of-the-art approaches on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. EBC-ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Method Venue MAE RMSE NAE
ChfL [33] CVPR’22 76.8 343.0 0.171
MAN [34] CVPR’22 76.5 323.0 0.170
CLTR [35] ECCV’22 74.4 333.8 0.165
CrowdHat [36] CVPR’23 68.7 296.9 0.181
STEERER [13] ICCV’23 63.7 309.8 0.133
PET [37] ICCV’23 74.4 328.5 -
APGCC [38] ECCV’24 71.7 284.4 0.215
Gramformer [39] AAAI’24 72.5 316.4 0.160
mPrompt [18] CVPR’24 62.1 293.5 -
CLIP-EBC [15] ICME’25 61.3 278.4 0.148
EBC-ZIP-B (ours) - 60.1 299.0 0.104
@@
TABLE V: Comparison of our method EBC-ZIP-B with the latest approaches on the test split of NWPU-Crowd. EBC-ZIP achieves both the lowest MAE and NAE on this large-scale benchmark.
IV-B EBC-ZIP-B Compared with State-of-the-Art
@@
We first compare our proposed model, EBC-ZIP-B (the backbone choice and computational complexity can be found in Table I), with recent state-of-the-art crowd counting methods on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table IV. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.81) and RMSE (75.04), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC [38]). On ShanghaiTech B, we also achieve the best performance with a MAE of 5.51 and RMSE of 8.63, significantly surpassing strong baselines such as CrowdHat [36] and STEERER [13]. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.46 MAE and 121.88 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.26 MAE and 64.84 RMSE, significantly outperforming the closest competitor, CLIP-EBC [15], which reports 36.6 MAE and 81.7 RMSE.
@@
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table V. Our EBC-ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104) among all methods, indicating superior overall accuracy and robustness. Although the RMSE (299.0) is not the lowest, it is comparable to other top-performing methods. Notably, RMSE tends to amplify the influence of outlier errors. Given our observation of annotation noise, such as duplicate or out-of-bound head labels in the training and validation splits, RMSE may not always reflect true performance. In contrast, MAE and NAE offer more stable and representative evaluations, making our results more reliable in practice. These results clearly demonstrate the effectiveness and generalizability of our method across various crowd densities and scenes.
@@
LSANet [25] 0.20 14.40 66.1 110.2 8.6 13.9 112.3 186.9
EBC-ZIP-P (ours) 0.81 6.46 71.18 109.6 8.23 12.62 96.29 161.82
EBC-ZIP-N (ours) 3.36 24.73 60.21 95.61 7.74 12.14 86.46 147.64
MobileCount [16] 3.40 16.49 89.4 146.0 9.0 15.4 131.1 222.6
LMSFFNet-XS [26] 4.58 39.28 85.8 139.9 9.2 15.1 112.8 201.6
MobleCount×$\times 1.25$ [16] 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
EBC-ZIP-T (ours) 10.53 61.39 59.07 90.55 6.67 9.90 76.02 129.40
MobleCount×$\times 2$ [16] 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
EBC-ZIP-S (ours) 33.60 242.43 57.20 93.07 5.88 9.37 73.32 125.09

### v2

@@ Introduction
The negative log-likelihood (NLL) loss of the zero-inflated Poisson distribution is given by:
$\mathcal{L}_{\mathrm{NLL}}=-\sum_{i=1}^{h}\sum_{j=1}^{w}\log P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ (7)
where $P(\boldsymbol{Y}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j})$ is the p.m.f. of the ZIP distribution given by Eq. (1), and $\boldsymbol{\lambda}$ and $\boldsymbol{\pi}$ are given by (4) and (6), respectively.
It is important to note that we do not supervise the $\boldsymbol{\pi}$ and $\boldsymbol{\lambda}$ branches separately. Since structural and sampling zeros are not explicitly annotated in the ground-truth, we treat them as latent factors and optimize both branches jointly via the ZIP negative log-likelihood loss Eq. (7). This formulation allows the model to implicitly learn to disentangle the two types of zeros based on the statistical patterns in the data, without requiring additional annotations.
@@
$\boldsymbol{Y}^{*}_{i,j}=\mathbb{E}[\boldsymbol{V}_{i,j}\mid\boldsymbol{\pi}_{i,j},\boldsymbol{\lambda}_{i,j}]=(1-\boldsymbol{\pi})\otimes\boldsymbol{\lambda},$ (8)
@@
ZIP Has a Tighter Risk Bound
Since Wang et al. established that DMCount enjoys a tighter risk bound than Gaussian-smoothed MSE-based losses, it suffices for us to show that the ZIP NLL loss defined in Eq. (7) admits an even tighter risk bound than DMCount. We formalize this result in the following theorem.
Theorem 1.
Assume the per-block counts follow a zero-inflated Poisson law with $\boldsymbol{\pi}>0$. Let $\mathcal{D}$ be the unknown ground-truth data distribution and $\mathcal{S}$ be the i.i.d. $K$ samples. Denote the hypothesis class of scalar regressors by $\mathcal{H}$ and the family of vector blockwise density predictors by =×$\mathcal{F}=\mathcal{H}^{\times hw}$. Let $f_{\mathcal{S}}^{\mathrm{NLL}}$ denotes an empirical risk minimizer of $\mathcal{L}_{\mathrm{NLL}}$ on $\mathcal{S}$, and $f_{\mathcal{D}}$ is the population minimizer (“Bayes optimal predictor”) for the same loss. Then, for any $\delta\in(0,1)$, with probability at least $1-\delta$ over the choice of the $K$ i.i.d. training examples $\mathcal{S}$, every $f\in\mathcal{F}$ satisfies
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{NLL}})$
$\leq 2\cdot hw\cdot C\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}),$ (12)
where $hw$ is the total number of blocks, $C$ is a dataset-specific constant, and $R_{\mathcal{S}}(\mathcal{H})$ is the empirical Rademacher complexity of the hypothesis class $\mathcal{H}$ on the sample set $\mathcal{S}$.
By comparison, generalization gap of the OT loss in DMCount (Wang et al. 2020a) is bounded by
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{OT}},\mathcal{L}_{\mathrm{OT}})-\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{OT}})$
$\leq$ $4\cdot(hw)^{2}\cdot C_{\infty}\cdot R_{\mathcal{S}}(\mathcal{H})+\mathcal{O}(\sqrt{1/K}).$ (13)
Hence, when $h$ or $w$ grows, the upper bound of ZIP (Eq. (12)) becomes tighter than that of OT (Eq. (13)). The proof of Theorem 1 can be found in the supplementary material.
Experiments
@@
Variants. To evaluate the scalability and adaptability of our proposed framework, we construct five model variants of ZIP with different backbones, including MobileNetV4 (Qin et al. 2024), MobileCLIP (Vasu et al. 2024) and OpenCLIP’s ConvNeXt-Base (Liu et al. 2022; Cherti et al. 2023) These variants are suitable for different application scenarios, ranging from mobile-friendly to high-performance settings. The specifications are summarized in Table 1. All variants share the same head design and differ only in backbone complexity. This modularity allows ZIP to scale across computational budgets.
@@
Setup. Our training configuration largely follows. We first preprocess all datasets so that the minimum edge length is no less than 448. We limit the maximum edge length of UCF-QNRF to be 2048 and that of NWPU-Crowd to be 3072. We apply a combination of data augmentation techniques, including random resized cropping, horizontal flipping, color jittering, and salt-and-pepper noise injection. For ShanghaiTech A and B, which consist of lower-resolution images, the crop size is set to $448\times 448$. For the higher-resolution datasets UCF-QNRF and NWPU-Crowd, we use a larger crop size of $672\times 672$. All models are optimized using the Adam optimizer with a weight decay of $41e-$ 4. The learning rate is linearly warmed up from $51e-$ 5 to $41e-$ 4 over the first 25 epochs, followed by cosine annealing with parameters $T_{0}=5$ and $T_{\mathrm{mult}}=2$.
@@
where $M$ denotes the total number of test images, $c_{m}$ is the ground-truth count for image $m$, and $c_{m}^{*}$ is the predicted count, computed via Eq. (9). The set $\mathcal{M}+\subseteq{1,2,\dots,M}$ includes only the indices for which $c_{m}>0$, and $\|\mathcal{M}_{+}\|$ denotes its cardinality.
Suffix Backbone Size (M) FLOPS (G)
-P (Pico) MobileNetV4-Small-0.5×0.5$\times$0.5 × 0.81 6.46
-N (Nano) MobileNetV4-Small 3.36 24.73
-T (Tiny) MobileNetV4-Medium 10.64 61.62
-S (Small) MobileCLIP-S1 33.72 242.68
-B (Base) CLIP-ConvNeXt-Base 105.60 800.99
Table 1: ZIP variants with corresponding backbones, parameter sizes, and computational complexity measured on $1920\times 1080$ resolution.
Component MAE RMSE NAE
MAE 6.69 10.45 5.52%
DMCount (Wang et al. 2020a) 6.62 9.83 5.30%
Poisson NLL 6.57 9.86 5.47%
$\mathcal{L}_{\mathrm{NLL}}$ (ours) 6.36 10.32 5.00%
Table 2: Comparison of $\mathcal{L}_{\mathrm{NLL}}$ (Eq. (7)) with alternative choices in the total loss function (Eq. (11)) on ShanghaiTech B, while $\mathcal{L}_{\mathrm{CE}}$ and $\mathcal{L}_{\mathrm{count}}$ remain unchanged. Results show that our ZIP NLL achieves the lowest MAE and NAE.
Method ShanghaiTech A ShanghaiTech B UCF-QNRF NWPU (Val)
@@
ChfL (Shu et al. 2022) 57.5 94.3 6.9 11.0 80.3 137.6 76.8 343.0
MAN (Lin et al. 2022) 56.8 90.3 - - 77.3 131.5 76.5 323.0
CLTR (Liang, Xu, and Bai 2022) 56.9 95.2 6.5 10.6 85.8 141.3 61.9 246.3
CrowdHat (Wu and Yang 2023) 51.2 81.9 5.7 9.4 75.1 126.7 - -
STEERER (Han et al. 2023) 54.5 86.9 5.8 8.5 74.3 128.3 54.3 238.3
PET (Liu et al. 2023) 49.3 78.8 6.2 9.7 79.5 144.3 58.5 238.0
APGCC (Chen et al. 2024) 48.8 76.7 5.6 8.7 80.1 133.6 - -
Gramformer (Lin et al. 2024) 54.7 87.1 - - 76.7 129.5 - -
mPrompt (Guo et al. 2024) 52.5 88.9 5.8 9.6 72.2 133.1 50.2 219.0
CLIP-EBC (Ma, Sanchez, and Guha 2024) 52.5 83.2 6.0 10.1 80.3 136.6 36.6 81.7
ZIP-B (ours) 47.81 75.04 5.51 8.63 69.46 121.88 28.26 64.84
@@
Table 3: Comparison of our model ZIP-B with state-of-the-art approaches on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Effectiveness of ZIP
In this subsection, we empirically validate the effectiveness of our proposed ZIP framework by comparing it against several strong baselines, including BL (Ma et al. 2019), DMCount (Wang et al. 2020a), and EBC (Ma, Sanchez, and Guha 2024). For a fair comparison, we adopt a consistent encoder-decoder architecture based on VGG19 (Simonyan and Zisserman 2015). Before proceeding to the comparison, we first determine the optimal value of the weighting parameter $\omega$ in Eq.(11), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig. 5. As illustrated, the model achieves the best MAE and RMSE when $\omega=0.75$, and the lowest NAE when $\omega=1.00$. In our subsequent comparisons, we fix $\omega=1.00$ to prioritize NAE, due to its robustness to varying crowd densities.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function (11) with several commonly used alternatives, including MAE, DMCount (Wang et al. 2020a), and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table 2. As shown, our ZIP NLL achieves the lowest MAE (6.36) and NAE (5.00%), and performs competitively in RMSE (10.32), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
Method MAE RMSE NAE
ChfL (Shu et al. 2022) 76.8 343.0 0.171
MAN (Lin et al. 2022) 76.5 323.0 0.170
CLTR (Liang, Xu, and Bai 2022) 74.4 333.8 0.165
CrowdHat (Wu and Yang 2023) 68.7 296.9 0.181
STEERER (Han et al. 2023) 63.7 309.8 0.133
PET (Liu et al. 2023) 74.4 328.5 -
APGCC (Chen et al. 2024) 71.7 284.4 0.215
Gramformer (Lin et al. 2024) 72.5 316.4 0.160
mPrompt (Guo et al. 2024) 62.1 293.5 -
CLIP-EBC (Ma, Sanchez, and Guha 2024) 61.3 278.4 0.148
ZIP-B (ours) 60.1 299.0 0.104
@@
Table 4: Comparison of our method ZIP-B with the latest approaches on the test split of NWPU-Crowd. ZIP achieves both the lowest MAE and NAE.
ZIP-B Compared with State-of-the-Art
@@
We first compare our proposed model, ZIP-B, with recent state-of-the-art crowd counting methods on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table 3. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.81) and RMSE (75.04), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC). On ShanghaiTech B, we also achieve the best performance with a MAE of 5.51 and RMSE of 8.63, significantly surpassing strong baselines such as CrowdHat and STEERER. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.46 MAE and 121.88 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.26 MAE and 64.84 RMSE, significantly outperforming the closest competitor, CLIP-EBC, which reports 36.6 MAE and 81.7 RMSE.
@@
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table 4. Our ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104) among all methods, indicating superior overall accuracy and robustness. Although the RMSE (299.0) is not the lowest, it is comparable to other top-performing methods. Notably, RMSE tends to amplify the influence of outlier errors. Given our observation of annotation noise, such as duplicate or out-of-bound head labels in the training and validation splits, RMSE may not always reflect true performance. In contrast, MAE and NAE offer more stable and representative evaluations, making our results more reliable in practice. These results clearly demonstrate the effectiveness and generalizability of our method across various crowd densities and scenes.
Comparison with Lightweight Models
@@
LSANet (Zhu et al. 2022) 0.20 14.40 66.1 110.2 8.6 13.9 112.3 186.9
ZIP-P (ours) 0.81 6.46 71.18 109.60 8.23 12.62 96.29 161.82
ZIP-N (ours) 3.36 24.73 58.86 94.63 7.74 12.14 86.46 147.64
MobileCount (Wang et al. 2020b) 3.40 16.49 89.4 146.0 9.0 15.4 131.1 222.6
LMSFFNet-XS (Yi et al. 2023) 4.58 39.28 85.8 139.9 9.2 15.1 112.8 201.6
MobleCount×$\times 1.25$ (Wang et al. 2020b) 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
ZIP-T (ours) 10.53 61.39 56.36 90.55 6.67 9.90 76.02 129.40
MobleCount×$\times 2$ (Wang et al. 2020b) 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
ZIP-S (ours) 33.60 242.43 55.17 88.99 5.83 9.21 73.32 125.09